OpenStudy (anonymous):

find the one-parameter family of solutions for (1+x^2)dy/dx=y

OpenStudy (shubhamsrg):

=> dx/(1+x^2) = dy/y integrate both sides..

OpenStudy (anonymous):

what do I do next?

OpenStudy (shubhamsrg):

integrate ,.thats the only thing you require to do ..

OpenStudy (shubhamsrg):

i mean,integrate both sides..

OpenStudy (anonymous):

I integrated both sides. There's nothing else to do? I get arctanx+c=lny+c

OpenStudy (shubhamsrg):

you may treat sum or difference of both constants as another constant C.. so i guess ans would be arctanx = lny + C

OpenStudy (shubhamsrg):

or you may want to solve for y from here..

OpenStudy (shubhamsrg):

for that,,let C = log k ,,where k = another constant => arctan x = lny + ln k => arctan x = ln(yk) =>e^(arctan x) /k = y or y = Ke^(arctan (x))

OpenStudy (shubhamsrg):

where K=1/k ..

OpenStudy (anonymous):

thank shubham!

OpenStudy (shubhamsrg):

your welcome,,glad to help..